The directions on a box of muffin mix calls for 1-1/4 cups of muffin mix and 1/2 cup of water. There is only 1 cup of muffin mix left in the box, how much water should be used?
Answer by rfer(16322)

1.25/0.5=1/x
1.25x=1*0.5
x=0.5/1.25
x=0.4 cups